# thanks: Deceasedthanks: Deceased

### v1

@@ II Data analysis and event discrimination
The data analysis has been performed with the REST-for-physics (Rare Event Searches Toolkit for Physics) framework [REST], and it can be fully replicated using the official version 2.4.0. The raw data in the form of voltage pulses are read from the detector and turned into a set of digitized waveforms, which are dubbed signal events in REST. The shape of the raw signals depends on parameters such as the shaping time, trigger delay or time duration of the acquisition window. Enough bins without signal are left at the beginning of the window, to be able to compute a baseline. After some processing and taking into account the detector and readout descriptions, these events turn into hits events with a given physical position and energy (and relative time). These hits can be grouped into track events if they fulfill a set of conditions that classify them as belonging to the same physical event, such as the maximum 3-dimensional distance (XYZ, or equivalently XY-time). This analysis chain is depicted in Fig. S4.
@@
The use of a stripped readout allows for topological analysis of the track events, which enables the definition of selection algorithms based on the topological shape of the events. As seen in Fig. S4, X-ray like events are small, symmetric, point like and consist of a single track, whereas background events have multiple tracks and are highly asymmetric (e.g. muon or alpha particle tracks). The process of Micromegas data cut optimization used in this work is explained in [MMcuts]. A further cut based on muons is applied. If an event in the Micromegas happens just after a muon signal is detected, the event is removed. Random coincidences are expected to be less than 0.5%, giving this cut a very high efficiency while simultaneously reducing the background level by a factor $\sim 1.7$
@@ III Efficiencies and detector response
Figure S5: Depiction of the total efficiency for each dataset and the separate components: detector (meaning gas absorption and window transmission) in blue, software in orange and telescope in green. The combined total efficiency is drawn in red. A horizontal line at 1 would represent 100% efficiency.
Monte Carlo simulations of the detector response have been performed with REST-for-physics [REST] and Geant4 [AGOSTINELLI2003250]. A flat X-ray flux between 0 and 12 keV from a point source placed in the center of the copper pipe, 100 mm away from the the detector window, was simulated.
@@
The final efficiency is about twice as high as the efficiency obtained in the last results [Anastassopoulos:2017ftl], and as seen in Fig.S7, the improvement is especially noticeable at low energies where the expected solar axion flux is much higher. This has a significant impact in the final result.
@@ IV Systematic effects
Magnetic field $B$ and magnet length $L$ The CAST magnet was run at a very stable current of 13 kA with negligible deviations from that value. The corresponding magnetic field is obtained by fitting data relating current and magnetic field (see e.g. [JVogelTesis]). A linear fit is performed and the error in the fit parameters is used to compute the uncertainty by error propagation taking correlation into account. The estimated uncertainty is $B=8.805\pm 0.034$ T. The nominal value for the magnet length provided by the magnet group is $L=(9.26\pm 0.05)$ m.
@@
Solar axion flux– Uncertainties in the theoretical solar axion flux and their effects in solar axion searches are studied in [2101.08789]. The flux can be expected to have a statistical fluctuation of $\sim 1.5\%$, and it is also model dependent. Helioseismological models, based on the internal structure and dynamics of the Sun using information from internal sound waves, yield Primakoff fluxes that are consistently $\sim 5\%$ higher than those predicted by photospheric models. As axions are expected to be produced in the inner 20% of the Sun, the former type of model is more appropriate, and thus we compute the model uncertainty by considering a flux 5% lower. The particular expression used in this work is derived in [Andriamonje:2007ew].
@@ V Expected sensitivity
To compute the expected sensitivity of the experiment one needs to simulate sets of candidates (i.e., X-ray like events during tracking) that follow the background distribution in a given readout area, for a given exposure time and energy interval. In a binned case, a representative dataset such as the Asimov dataset [Cowan_2011] provides an efficient solution to compute an expected limit. However, in an unbinned case a set of simulations has been performed instead, which also allows to study the possible variation due to statistical fluctuations. The expected sensitivity of the experiment can be simulated as follows:

### v3

@@ II Data analysis and event discrimination
The data analysis has been performed with the REST-for-physics (Rare Event Searches Toolkit for Physics) framework [2], and it can be fully replicated using the official version 2.4.0. The raw data in the form of voltage pulses are read from the detector and turned into a set of digitized waveforms, which are dubbed signal events in REST. The shape of the raw signals depends on parameters such as the shaping time, trigger delay or time duration of the acquisition window. Enough bins without signal are left at the beginning of the window, to be able to compute a baseline. After some processing and taking into account the detector and readout descriptions, these events turn into hits events with a given physical position and energy (and relative time). These hits can be grouped into track events if they fulfill a set of conditions that classify them as belonging to the same physical event, such as the maximum 3-dimensional distance (XYZ, or equivalently XY-time). This analysis chain is depicted in Fig. S4.
@@
The use of a stripped readout allows for topological analysis of the track events, which enables the definition of selection algorithms based on the topological shape of the events. As seen in Fig. S4, X-ray like events are small, symmetric, point like and consist of a single track, whereas background events have multiple tracks and are highly asymmetric (e.g. muon or alpha particle tracks). The process of Micromegas data cut optimization used in this work is explained in [1]. A further cut based on muons is applied. If an event in the Micromegas happens just after a muon signal is detected, the event is removed. Random coincidences are expected to be less than 0.5%, giving this cut a very high efficiency while simultaneously reducing the background level by a factor $\sim 1.7$
@@ III Efficiencies and detector response
Figure S5: Depiction of the total efficiency for each dataset and the separate components: detector (meaning gas absorption and window transmission) in blue, software in orange and telescope in green. The energy threshold is taken into account as part of the software efficiency, and is represented as a sharp drop at 0.5 keV for Ar and 1 keV for Xe. The combined total efficiency is drawn in red. A horizontal line at 1 would represent 100% efficiency.
Monte Carlo simulations of the detector response have been performed with REST-for-physics [2] and Geant4 [3]. A flat X-ray flux between 0 and 12 keV from a point source placed in the center of the copper pipe, 100 mm away from the the detector window, was simulated.
@@
The final efficiency is about twice as high as the efficiency obtained in the last results [4], and as seen in Fig.S7, the improvement is especially noticeable at low energies where the expected solar axion flux is much higher. This has a significant impact in the final result.
@@ IV Systematic effects
Magnetic field $B$ and magnet length $L$ The CAST magnet was run at a very stable current of 13 kA with negligible deviations from that value. The corresponding magnetic field is obtained by fitting data relating current and magnetic field (see e.g. [5]). A linear fit is performed and the error in the fit parameters is used to compute the uncertainty by error propagation taking correlation into account. The estimated uncertainty is $B=8.805\pm 0.034$ T. The nominal value for the magnet length provided by the magnet group is $L=(9.26\pm 0.05)$ m. A 0.93346 m long haloscope cavity was installed inside the bore, reducing the effective magnetic field length.
@@
Solar axion flux– Uncertainties in the theoretical solar axion flux and their effects in solar axion searches are studied in [6]. The flux can be expected to have a statistical fluctuation of $\sim 1.5\%$, and it is also model dependent. Helioseismological models, based on the internal structure and dynamics of the Sun using information from internal sound waves, yield Primakoff fluxes that are consistently $\sim 5\%$ higher than those predicted by photospheric models. As axions are expected to be produced in the inner 20% of the Sun, the former type of model is more appropriate, and thus we compute the model uncertainty by considering a flux 5% lower. The particular expression used in this work is derived in [7].
@@ V Expected sensitivity
To compute the expected sensitivity of the experiment one needs to simulate sets of candidates (i.e., X-ray like events during tracking) that follow the background distribution in a given readout area, for a given exposure time and energy interval. In a binned case, a representative dataset such as the Asimov dataset [8] provides an efficient solution to compute an expected limit. However, in an unbinned case a set of simulations has been performed instead, which also allows to study the possible variation due to statistical fluctuations. The expected sensitivity of the experiment can be simulated as follows: